Directions: Calculate the mean, median, mode, range, variance, and standard deviation (SD) for each set of data. Please show your work on a separate sheet of paper and submit it along with this worksheet. Make sure your name is on the separate sheet. All answers must be written on this sheet.

1. Data Set: 1, 3, 1, 5, 7, 2, 4, 1, 3, 6, 2, 5, 2, 6, 8, 8, 2, 1, and 3

         

Mean = _____         Median=_____        Mode=_____           Range=_____

Variance=_____      SD=________

2. Data Set: 11, 14, 87, 34, 26, 15, 26, 11, 24, 27, and 15

Mean = _____         Median=_____        Mode=_____           Range=_____

Variance=__________       SD=_______

3. Data Set: 131, 252, 313, 131, and 145

Mean = _____         Median=_____        Mode=_____           Range=_____

Variance=__________       SD=_______

4. Data Set: 1,273, 3,987, 2,763, 2,758, 3,000, 2,964, and 1,997

Mean = _____         Median=_____        Mode=_____           Range=_____

Variance=__________       SD=_______

5. Data Set: 16, 27, 131, 1,237, 3,987, 14, 87, 32, 31, 87, 26, 15, and 11

Mean = _____         Median=_____        Mode=_____           Range=_____

Variance=__________       SD=_______

6. A study is conducted to see how effective aspirin is in reducing temperature in children. A sample of 7 children suffering from influenza had their temperatures taken immediately before and 1 hour after administration of aspirin. The results are given below. The data follow.

Patient

Temperature Before

Temperature After

1

99.7

99

2

101.7

100.9

3

101.7

100.8

4

100.7

99

5

101.7

98.4

6

103.7

103.5

7

99.7

97

Note: Children under 17 should not be given aspirin unless advised by a physician, due to the possibility of Reyes Syndrome, a disease affecting the liver and brain.

  1. Calculate the mean, median, mode, and range for each set of values (before and after)

Before:                                                           After:

Mean: _______                                               Mean: _________

Median: ______                                              Median: _______

Mode: _______                                               Mode: _________

Range: _______                                              Range: _________

  1. Calculate the variance and standard deviation for each set of values (before and after)

Before:                                                           After:

Variance: _______                                          Variance: _______

SD: _________                                               SD: _______
Answer #1

1. Data set - 1, 3, 1, 5, 7, 2, 4, 1, 3, 6, 2, 5, 2, 6, 8, 8, 2, 1, and 3

Mean = sum of the observation / number of observations

= 1+3+1+5+7+2+4+1+3+6+2+5+2+6+8+8+2+1+3 / 19

= 71/19 = 3.73

Median - arrange the set in ascending order

1,1,1,1,2,2,2,2,3,3,3,4,5,5,6,6,7,8,8

Since the number of observation is 19 which is odd.

So the number coming in middle is the median = 3

Mode - the observation which repeat often is the mode = 1 and 2

Range is the difference between largest observation and smallest observation = 8-1 = 7

2. Data Set: 11, 14, 87, 34, 26, 15, 26, 11, 24, 27, and 15

Mean = sum of the observation / number of observations

= 11+14+87+34+26+15+26+11+24+27+15/ 11

= 290/11

= 26.3

Median - arrange the set in ascending order

11,11,14,15,15,24,26,26,27,34,87

Since the number of observation is 11 which is odd.

So the number coming in middle is the median = 24

Mode - the observation which repeat often is the mode = 11,15 and 26

Range is the difference between largest observation and smallest observation = 87- 11= 76

3. Data Set: 131, 252, 313, 131, and 145

Mean = 131+252+313+131+145 / 5 = 856

Median = arrange in ascending order

= 131,131,145,252,313

= 145

Mode - is the one repeats often

= 131

Range = 313-131= 182

4. Data Set: 1,273, 3,987, 2,763, 2,758, 3,000, 2,964, and 1,997

Mean = 1273+3987+2763+2758+3000+2694+1997 / 7=

Median = 1273, 1997, 2758,2763, 2964, 3000,3987

= 2763

Mode = Nil

Range = 3987- 1273 =2714